OpenStudy (anonymous):

What is the expression in factored form? 16x2 - 25

OpenStudy (anonymous):

the problem has two terms in it, so if it can be factored we'll have to use the difference of squares. so use this we have to see if both terms are squares, are they?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

A. (4x + 5)(4x - 5) B. (4x - 5)2 C. (4x + 5)(-4x - 5) D. (-4x + 5)(4x - 5)

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

this one is a bit tricky, ill show you 16x^2 = 4^2 * x^2 = (4x)^2 so we know the first term is a square, what is the root of 25?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

great so we can rewrite our equation as (4x)^2 - 5^2 to apply the 'difference of squares' a^2 - b^2 = (a-b)(a+b) do you see how 4x is like the 'a' and 5 is like the 'b'?

OpenStudy (anonymous):

it would have to be a. if you take and factor out the answer choices, (4x+5)(4x-5)= 16x^2-20x+20x-25. the 20x's cancel which give you 16x^2-25